# Regularized Black Hole Solution from a New String Cloud Source

### v1

@@ VIII Conclusions
In this work, we have constructed a new family of regular black hole solutions supported by the Letelier-Alencar string cloud and regularized through a rational Dagum-type distribution. This regulator successfully smooths the matter profile and ensures the finiteness of curvature invariants, overcoming the limitations of the standard exponential regularization adopted in previous extensions of the Letelier model. The resulting geometry interpolates between a string–cloud–dominated regime at large distances and an anti–de Sitter vacuum core at the origin, characterized by an effective curvature radius determined by the string parameter $a$ and the regularization scale $r_{0}$.
We analyzed in detail the corresponding energy density profile and showed how the interplay between the string cloud strength and the regularization scale controls the transition between the AdS-like inner region and the asymptotically flat exterior governed by the string cloud. The full set of energy conditions was investigated, allowing us to identify the domains where the null, weak, dominant, and strong conditions are satisfied or violated, in agreement with the expected structure of regular black holes with a smooth center.
A complete thermodynamic analysis was performed. We derived the mass, Hawking temperature, entropy, and heat capacity of the solution, demonstrating in particular that while both $a$ and $r_{0}$ influence the temperature and mass distribution, the black hole entropy depends solely on the regularization scale. This indicates that the microscopic degrees of freedom associated with the entropy are governed by the regular core rather than by the external string field. We also studied the non-extensive thermodynamics of the system through the Rényi entropy formalism, showing that the deformation parameter $\lambda$ smooths out the phase structure and eliminates the standard phase transition, yielding a single thermodynamically stable branch. These results were corroborated by the topological thermodynamics approach, which identifies black hole states as topological defects and confirms the absence of phase transitions in the non-extensive regime.
@@
In summary, the regularized string cloud black hole presented here constitutes a consistent and physically motivated extension of recent developments in regular geometries, combining mathematical regularity, well-defined thermodynamic behavior, and phenomenological viability.

### v2

@@ VIII Conclusions
In this work, we constructed a new family of regular black hole solutions supported by the Letelier–Alencar string cloud and regularized it through a rational Dagum-type distribution, a regulator that has also been used in the literature to generate regular geometries sourced by nonlinear electrodynamics. The employed regularization successfully smooths the matter profile and ensures the finiteness of curvature invariants, overcoming the limitations of the standard exponential regularization adopted in previous extensions of the Letelier model. The resulting geometry interpolates between a string–cloud–dominated regime at large distances and an anti–de Sitter vacuum core at the origin, characterized by an effective curvature radius determined by the string parameter $a$ and the regularization scale $r_{0}$. In addition, we have shown that our regularized string–cloud black hole and the Frolov regular black hole solution exhibit the same near-origin behavior up to $\mathcal{O}(r^{4})$, as demonstrated by the matching leading terms in the expansions of the metric function and curvature invariants. Although this local agreement confirms that both geometries share the same regular core structure – AdS-type – it does not imply that their physical sources are equivalent.
We analyzed in detail the matter source associated with our solution, showing from the time component of the field equations that the resulting energy density profile captures how the interplay between the string cloud strength and the regularization scale governs the transition from the AdS-like inner region to the asymptotically flat exterior dominated by the string cloud. By examining the remaining components of Einstein’s equations, we evaluated the full set of energy conditions and identified the regions where the null, weak, dominant, and strong conditions are satisfied or violated, in accordance with the expected behavior of regular black holes possessing a smooth central core.
A complete thermodynamic analysis was performed. We derived the mass, Hawking temperature, entropy, and heat capacity of the solution, demonstrating in particular that while both $a$ and $r_{0}$ influence the temperature and mass distribution, the black hole entropy, where the Hawking-Bekenstein area law is modified by logarithmic and inverse-radius corrections introduced by the regularization scale $r_{0}$, depends solely on this scale. This indicates that the microscopic degrees of freedom associated with the entropy are governed by the regular core rather than by the external string field. We also studied the non-extensive thermodynamics of the system through the Rényi entropy formalism, showing that the deformation parameter $\lambda$ smooths out the phase structure and eliminates the standard phase transition, yielding a single thermodynamically stable branch. These results were corroborated by the topological thermodynamics approach, which identifies black hole states as topological defects and confirms the absence of phase transitions in the non-extensive regime.
@@
In summary, the regularized string cloud black hole presented here constitutes a consistent and physically motivated extension of recent developments in regular geometries, combining mathematical regularity, well-defined thermodynamic behavior, and phenomenological viability. Future investigations may explore broader theoretical settings, possible generalizations of the matter sector as well as refinements and extensions of the phenomenological analysis, offering additional insight into the role of regular cores and extended matter fields in gravitational physics.